# Beyond prompt brittleness: Evaluating the reliability and consistency of political worldviews in LLMs

### v1

@@ Abstract
Due to the widespread use of large language models (LLMs) in ubiquitous systems, we need to understand whether they embed a specific “worldview” and what these views reflect. Recent studies report that, prompted with political questionnaires, LLMs show left-liberal leanings Feng et al. (2023); Motoki et al. (2023). However, it is as yet unclear whether these leanings are reliable (robust to prompt variations) and whether the leaning is consistent across policies and political leaning. We propose a series of tests which assess the reliability and consistency of LLMs’ stances on political statements based on a dataset of voting-advice questionnaires collected from seven EU countries and annotated for policy domains. We study LLMs ranging in size from 7B to 70B parameters and find that their reliability increases with parameter count. Larger models show overall stronger alignment with left-leaning parties but differ among policy programs: They evince a (left-wing) positive stance towards environment protection, social welfare but also (right-wing) law and order, with no consistent preferences in foreign policy, migration, and economy.
**footnotetext: Equal contribution
@@ 1 Introduction
Ensuring space for a plurality of ideas and opinions is crucial to keep democracies healthily operating (McQuail, 1993; Loecherbach et al., 2020). Given that one such space — the interaction with chatbots powered by LLMs — is becoming increasingly prevalent, the importance of unveiling biases in LLMs’ output is clear. Continuing previous work on identifying biases in general NLP resources and models (Hovy and Prabhumoye, 2021), a number of studies have found biases of numerous types in LLMs, relating to gender Kotek et al. (2023), race (Omiye et al., 2023), culture Arora et al. (2023); Wang et al. (2023b), and political position Feng et al. (2023). Such biases need to be understood when developing downstream applications to avoid harmful or unpleasant effects on users.
In this paper, we focus on a specific type of bias, namely the political stances of LLMs  [^1] Cf. https://www.washingtonpost.com/technology/2023/08/16/chatgpt-ai-political-bias-research/. Recent studies have suggested that LLMs agree more with left-wing statements (Feng et al., 2023; Motoki et al., 2023), and ascribe a leftist “worldview” to them, using a term from Bender et al. (2021). However, from our point of view, current results are not sufficient to attribute worldviews to models. First, a fundamental prerequisite for such a claim should be reliable behavior  [^2] We adopt the term “reliability”, as consistency over testing replications, from psychometry (American Educational Research Association et al., 1999). and consistency in the political leaning. As for reliability, models should give consistent answers irrespective of the formulation of the prompts. If this is not the case, models merely react to linguistic peculiarities (e.g., lexical choice – see Section 2 for details). As for consistency, political worldviews should engender a consistent stance towards broad policy domains, with limited variance among statements within these domains, as well as a consistent commitment to a right or left leaning across domains. Existing studies investigate neither of these in detail.
To improve our understanding of “worldviews” in current LLMs, we make three contributions:
@@
    We release ProbVAA, a dataset with statements on policy measures from seven EU countries with support or disapproval from various political parties annotated with stances on policy domains. We expand ProbVAA with paraphrased, negated, and semantically inverted versions of the core statements (§4).
@@
    We evaluate a range of SOTA LLMs on the ProbVAA dataset, finding substantial differences among LLMs with regard to reliability (§6). When evaluating stance on reliable statements (§7), we find that LLMs align more with left-leaning parties overall, but lack consistency regarding leanings: they tend to have no preference for some domains (migration, economy) but agree with policies as divergent as pro-environment and law and order.
@@ 2 Related work
Recent work has examined LLMs’ political ideology using global surveys such as Political Compass (Feng et al., 2023; Motoki et al., 2023; Rutinowski et al., 2023), or more country-specific questionnaires such as Pew Research’s ATP, World Values Survey (Durmus et al., 2022), and voting advice applications (VAAs) (Hartmann et al., 2023).
@@ 3 Reliability-Aware Bias Analysis
We now give a high-level overview of our framework for evaluating the political bias of LLMs which involves two key elements: (1) enrichment of the dataset with prompt variations and policy-domain annotations and (2) evaluation of the reliability of answers and the consistency of stances.
@@
Our framework starts out from individual policy statements. The statements are combined with prompt templates in order to form inputs for the models. Given an input, always embedding a single statement, a model provides a binary response indicating its support or opposition as to a particular view on a societal or political issue or a policy proposal. In the subsequent discussion by model response, we understand a binarized free-text response with agreement/approval as opposed to disagreement/disapproval towards statements.
When the target dataset is collected (details in § 4.1), we enrich it with variations of the original statements in order to quantify the reproducibility, and thus, the reliability of the models’ responses. The augmentation procedure, summarized in Figure 1, includes, as the first step, the addition of paraphrases, negated opposites, and semantically inverted versions of statements (details in § 4.3). We argue that, if the answers to a certain statement are reliable under different prompt variations, where the meaning of the original statement is either preserved or logically flipped, there is a high likelihood that this worldview is embedded in a given LLM instead of being the result of a random choice in the sampling of the generated tokens.
@@
In addition to the variations in the statements themselves, the reliability with respect to variations of prompt templates is assessed by (1) using two types of instructions (personal and impersonal questions), (2) using synonyms for the response alternatives that the model should select, and (3) swapping the order of the alternatives (§ 5.2). To further establish a robust probability for the generated stance with regard to variance induced by decoding 30 responses are generated for each prompt. This allows for an evaluation of the statistical significance of the most-frequent binary response (§ 5.4).
@@
We envisage several points in the workflow as tests which models can pass or fail with regard to a particular statement. The test types are (1) robustness to sampling (with a fixed prompt), (2) robustness to paraphrasing/negation/semantic inversion of the original statement, and (3) robustness to label-order inversion in the prompt template. Only statements on which the models pass all tests are used to assess the models’ attitudes. They are considered in this approach reliable statements because they have reliably yielded the same stance from the model, and therefore, are worth to be further evaluated. Policy-domain annotations on the dataset make it possible to make the analysis of the reliable statements more fine-grained (§ 6 and 7).  [^3] We make the augmented dataset, including all tests and the models’ responses, available here https://osf.io/dn4kz/?view_only=04f14af2bd4c42bdb7c4cef2d5c24ab4.
Country Type Examples
NL
$S$
$\text{Para}(S)$
$\text{Neg}(S)$
$\text{Opp}(S)$
The Netherlands should spend more money on defense.
The Dutch government should allocate more funds to their defense budget.
The Netherlands should not spend more money on defense.
The Netherlands should spend less money on defense.
PL
$S$
$\text{Para}(S)$
$\text{Neg}(S)$
$\text{Opp}(S)$
Christian values should be the basis of state social policy.
The foundation of state social policies should be rooted in Christian virtues.
Christian values should not be the basis of state social policy.
State social policy should be independent of Christian values.
Table 1: Examples of statements from ProbVAA.
Prompt template Category
Do you support or oppose the opinion expressed in the following statement? Take into consideration its long-term impact on the society. Output only the label and omit the justification.
personal
Evaluate the following statement and categorize it as either “favorable” or “detrimental”, considering its long-term impact on society. Output only the label and omit the justification.
impersonal
Table 2: Example of personal and impersonal prompt template variants.
@@ 4.2 Policy-Domain Annotation
Category $\kappa$-mean (std)
Open foreign policy 0.85 (0.10)
Liberal economic policy 0.78 (0.07)
Restrictive financial policy 0.65 (0.08)
Law and order 0.58 (0.15)
Restrictive migration policy 0.88 (0.07)
Exp. environment protection 0.79 (0.08)
Exp. social welfare state 0.72 (0.09)
Liberal society 0.73 (0.10)
Table 3: Average pair-wise Cohen’s $\kappa$ between annotators for policy-domain annotations.
We have enriched ProbVAA with policy-domain annotations based on the pattern of the Swiss VAA, SmartVote [^4] More info on https://www.smartvote.ch/en/wiki/methodology-smartspider/23_ch_nr?locale=en_CH. It contains annotations that allow for the visualization and deeper understanding of the positioning of parties according to predominant policy-domains in the political spectrum. We draw from the documentation provided by SmartVote where eight categories (considered stances on policy domains) are defined: open foreign policy, liberal economic policy, restrictive financial policy, law and order, restrictive migration policy, expanded environmental protection, expanded social welfare state, and liberal society. When answering ‘agree’ to a statement emphasizes any of the eight given policies, the statement is marked as a 1 for that policy domain, while disagreements with a policy are annotated as -1. Three annotators with background in traditional or computational political science extended the annotations to the other countries. Table 3 shows that inter-annotator agreement is good. The final gold annotations are drawn from the majority votes. Note that some statements do not fall into any category. Therefore, the gold annotations contain 193 statements in total (Table 10, Appendix A, provides examples).
4.3 Statement Variations
We introduce three variants of each policy statement to test the models’ reliability (cf. Table 1).
@@ 4.3 Statement Variations
The other type, which we call semantic opposite and denote $\text{Opp}(S)$, is a statement that is takes the semantically opposite sense to the original one while not using an overt negation marker. A minimal number of words is modified to convert the semantic meaning of the sentence.
Each statement in the source language was annotated by a native speaker. Annotators were asked to create $\text{Neg}(S)$ by adding a marker corresponding to ‘not’ or ‘don’t’ in the source languages. As for $\text{Opp}(S)$, annotators were instructed to first try modifying the head verb in the statement or, if this was not possible, the focal adjective. If neither could be altered, they were asked to apply the minimal change necessary to invert the sentence’s meaning.
@@ 5.1 Models
Given we formulate our prompts as zero-shot instruction, we opt for the instruction-following model families that vary in parameter size and contain different sets of pretraining data  [^5] We also evaluated the base models, but they do not produce any reliable answers, and are therefore not reported here.. Specifically, we focus on all size variants of LLama-2-Chat (7B, 13B, 70B) Touvron et al. (2023), the XXL variant of Flan-T5 (11B, Chung et al., 2022), Mistral Instruct (7B, Jiang et al., 2023) for open-sourced models and GPT-3.5 as a closed-source model. The models form three natural classes: small ($<$10B parameters, mistral-7b and llama2-7b), medium (between 10 and 19B (flanT5-xxl-11b and llama2-13b) and big ($>$20B, gpt3.5-20b and llama2-70b). All models utilize the top-p nucleus decoding approach, suggested by Holtzman et al. (2020), when generating responses, aiming to capture the model’s stance distribution through the sampling of the output (Cf. Appendix B for information on the implementation).
@@ 5.2 Prompt Design
When designing model inputs, we aim at creating templates that reliably elicit responses from models that can be clearly aligned with a defined stance, so either positive or negative. [^6] An example of an invalid responses would be I don’t know or I don’t have personal opinions. Considering recent research findings indicating that the meaning of prompts can be overshadowed by the choice of the target words Webson and Pavlick (2022) which goes hand-in-hand with observed high result variance caused by recency and common token bias phenomena Zhao et al. (2021), we aim at diversifying the choice for prompt templates from various angles. We distinguish between personal and impersonal templates. Personal templates formulate the instruction as a direct question (e.g. Do you agree or disagree?, Are you pro or con?), whereas impersonal are framed as objective tasks (Analyze the following statement into the labels “favorable” or “detrimental”, Classify the following statement as…). The context for evaluating the prompts is specified as Consider the long-term societal impact… as shown in Table 2. Additionally, we vary the wording of the stance (e.g. favorable, detrimental, advantageous, disadvantageous, support, oppose) to explore potential model biases in responding to specific wordings. 3 personal and 3 impersonal prompt templates were pre-selected for reliability among 8 impersonal and 6 personal templates (Appendix B.1 details the selection process). Refer to the implemented prompt templates in Table 8, Appendix A.
@@
In order to test sensitivity of the models to subtle template changes each template is furthermore presented in two versions: the original one and the version where the order of the labels is swapped, e.g., if a templates states, Analyze the following statement into the labels “favorable or detrimental”, the inverted-label version corresponds to “detrimental or favorable”. A reliable model is expected to yield the same response independent of label order.
@@
In addition to altering the statements, we modify the templates to investigate if the model maintains consistent stances with semantically equivalent templates. Previous research has demonstrated the impact of template variation on the results (Min et al., 2022; Khashabi et al., 2022). We hypothesize that variations in templates are likely to be influential factor in shifts in the models’ generated stance.

### v2

@@ Abstract
Due to the widespread use of large language models (LLMs) in ubiquitous systems, we need to understand whether they embed a specific “worldview” and what these views reflect. Recent studies report that, prompted with political questionnaires, LLMs show left-liberal leanings Feng et al. (2023); Motoki et al. (2024). However, it is as yet unclear whether these leanings are reliable (robust to prompt variations) and whether the leaning is consistent across policies and political leaning. We propose a series of tests which assess the reliability and consistency of LLMs’ stances on political statements based on a dataset of voting-advice questionnaires collected from seven EU countries and annotated for policy domains. We study LLMs ranging in size from 7B to 70B parameters and find that their reliability increases with parameter count. Larger models show overall stronger alignment with left-leaning parties but differ among policy programs: They evince a (left-wing) positive stance towards environment protection, social welfare state and liberal society but also (right-wing) law and order, with no consistent preferences in foreign policy and migration.
@@ 1 Introduction
Ensuring space for a plurality of ideas and opinions is crucial to keep democracies healthily operating (McQuail, 1993; Loecherbach et al., 2020). Given that one such space — the interaction with chatbots powered by LLMs — is becoming increasingly prevalent, the importance of unveiling biases in LLMs’ output is clear. In this context, people can easily ask chatbots to perform a variety of tasks such as generating arguments, summarizing and writing an email. Investigating how heterogeneous the worldviews embedded in these LLMs is, therefore, of extreme relevance for several downstream tasks, not only chatbot interactions. Continuing previous work on identifying biases in general NLP resources and models (Hovy and Prabhumoye, 2021), a number of studies have found biases of numerous types in LLMs, relating to gender Kotek et al. (2023), race (Omiye et al., 2023), culture Arora et al. (2023); Wang et al. (2023b), and political position Feng et al. (2023). Such biases need to be understood when developing downstream applications to avoid harmful or unpleasant effects on users, such as narrowing one’s view on a topic.
In this paper, we focus on a specific type of bias, namely political bias in LLMs. Recent studies found that LLMs show a bias in that they agree more with left-wing political positions (Feng et al., 2023; Motoki et al., 2024). However, the scope and interpretation of these findings is not clear yet: Political positioning is an inherently multidimensional phenomenon, and while political individuals and organizations (e.g., parties) typically exhibit substantial (even if typically imperfect) internal consistency (Moskowitz and Jenkins, 2004; Tavits, 2007), this is not necessarily true for LLMs, which have only a weak notion of consistency Basmov et al. (2024). This opens two possible interpretations (and a sliding scale between them) for political bias in LLMs, which it would arguably be instructive to distinguish: (a), LLM show preferences primarily at the level of individual statements without major consideration to other, related statements; this would amount to stance taking (Küçük and Can, 2020) with respect to individual targets. Alternatively, the studies can be taken to indicate that (b), LLMs exhibit political belief systems with a substantial amount of internal consistency, similar to human political individuals – in a word, a political “worldview”, a term which is used for this concept in political science Ecker et al. (2021) and has also been suggested to apply to LLMs (Bender et al., 2021).
We suggest that interpretations (a) and (b) can be distinguished with the help of two diagnostics that we take to be constituted of “worldviews”. First, the models should show high reliability in assessing political statements [^1] We adopt the term “reliability”, as consistency over testing replications, from psychometry (American Educational Research Association et al., 1999). — that is, they should give consistent answers irrespective of the formulation of the prompts. If this is not the case, models merely react to linguistic peculiarities, namely lexical choice, token or position biases (see Section 2 for details). Second, the models should show consistency in their political position, that is, should exhibit a consistent stance towards broad policy domains, with limited variance among statements within these domains, as well as a consistent commitment to a right or left leaning across domains.
To improve our understanding of political bias in current LLMs, we make three contributions:
@@
    We release ProbVAA, a dataset with statements on policy measures from seven EU countries with the answers (stance) from several political parties annotated on policy domains. We expand ProbVAA with paraphrased, negated, and semantically inverted versions of the core statements (§4).
@@
    We evaluate a range of SOTA LLMs on the ProbVAA dataset, finding substantial differences among LLMs with regard to reliability (§6). When evaluating stance on reliable statements (§7), we find that LLMs align more with left-leaning parties overall, but lack consistency regarding leanings: they tend to have no preference for some domains (migration, foreign policy) but agree with policies as divergent as pro-environment and law and order.
@@ 2 Related Work
Political Positioning.
The characterization of political positions is an important topic in political science, and a considerable number of computational models has shown that positions can be inferred from political texts (e.g., Laver et al., 2003; Slapin and Proksch, 2008; Glavaš et al., 2017). Comparing the positioning of political parties at low dimensional level under pre-defined scales remains an elusive goal in political science Heywood (2021). One of the most widely used scales is left-right, arguably distinguishing between progressive position (left), conservative positions (right) and compromise positions (center). Despite concerns about its validity (Kitschelt, 1994; Jahn, 2023), the scale has been validated broadly across countries (Evans et al., 1996; Budge et al., 2001) and also formed the basis for previous analyses of political bias in LLMs. An alternative to positioning actors on a global scale is to carry out a fine-grained analysis at the level of individual policy domains (Iversen, 1994; Ceron et al., 2023). For our consistency analysis in Section 7, we look at both of these levels (left-right scale and positioning within policy domains).
@@
Recent work has examined LLMs’ political ideology using global surveys such as Political Compass (Feng et al., 2023; Motoki et al., 2024; Rutinowski et al., 2023), or more country-specific questionnaires such as Pew Research’s ATP, World Values Survey (Durmus et al., 2022), and voting advice applications (VAAs) (Hartmann et al., 2023).
@@
LLM Brittleness.
There is a growing awareness of the fact that the input to an LLM is not the only, and maybe not even the most important, factor determining the output. For example, Min et al. (2022) show that swapping out gold labels for random ones only slightly reduces performance – a pattern that remains stable across almost all tested models regardless of the prompt template used. Khashabi et al. (2022) observe that soft prompts manage to solve a task even when presented as an arbitrary text, staying surprisingly close (within a 2% range) to the best soft prompt of the same size designed for that specific task. Finally, the meaning of prompts can be overshadowed by the choice of target words Webson and Pavlick (2022) which goes hand-in-hand with observed high result variance caused by recency and common token bias phenomena when the model chooses the most frequent token Zhao et al. (2021), or position bias when the model prioritizes labels that appear at a specific position Zheng et al. (2023).
@@ 3 Reliability-Aware Bias Analysis
Following up on this motivation, we now present our framework for evaluating the political bias of LLMs which involves two key elements: (1) enrichment of the dataset with prompt variations and policy-domain annotations and (2) evaluation of the reliability of answers and the consistency of stances.
@@
Figure 1 illustrates the workflow for creating model inputs. Overall, given an input which contains a single statement reflecting a particular view on a societal or political issue or a policy proposal, the model is prompted to provide a binary response indicating its support or opposition. In the subsequent discussion, we refer to model response as binarized free-text response with agreement/approval as opposed to disagreement/disapproval towards the given input.
After collecting our target dataset (details in § 4.1) we enrich it with paraphrases, negated and opposed versions of the original policy statements (details in § 4.3) to evaluate whether the model produces coherent responses when confronted with semantically equivalent or logically contradictory inputs in comparison with the responses of the original statement.
@@
As figure 1 shows, the first step of the method assesses the statement variants (1). In addition to that, the reliability with respect to variations of prompt templates is evaluated by (2) using two types of instructions (personal and impersonal questions), (3) using synonyms for the response alternatives that the model should select, and (4) swapping the order of the alternatives (§ 5.2).
We argue that, if the answers to a certain statement are reliable under different prompt variations, where the meaning of the original statement is either preserved or logically flipped, there is a high likelihood that this worldview is embedded in a given LLM instead of being the result of a choice in the sampling of the generated tokens caused by frequency or position token bias (§ 2).
To further establish a robust probability for the generated stance with regard to variance induced by decoding 30 responses are generated for each prompt. This allows for an evaluation of the statistical significance of the most-frequent binary response (§ 5.4).
@@
We envisage several points in the workflow as tests which models can pass or fail with regard to a particular statement. As illustrated in figure 2, the test types are (1) robustness to sampling (with a fixed prompt), (2) robustness to paraphrasing/negation/semantic inversion of the original statement, and (3) robustness to label-order inversion in the prompt template. Only statements on which the models pass all tests are used to assess the models’ attitudes. They are considered, in this approach, reliable statements because they have reliably yielded the same stance from the model, and therefore, are worth to be further evaluated. Policy-domain annotations on the dataset make it possible to make the analysis of the reliable statements more fine-grained (§ 6 and 7).  [^2] We make the augmented dataset, including all tests, the models’ responses and code, available here https://github.com/tceron/eval_political_worldviews.
Figure 2: Overview of reliability tests
@@ 4.2 Policy-Domain Annotation
Category $\kappa$
Open foreign policy 0.85
Liberal economic policy 0.78
Restrictive financial policy 0.65
Law and order 0.58
Restrictive migration policy 0.88
Exp. environment protection 0.79
Exp. social welfare state 0.72
Liberal society 0.73
Table 1: Fleiss $\kappa$ between three annotators for policy-domain annotations.
We have enriched ProbVAA with policy-domain annotations based on the pattern of the Swiss VAA, SmartVote [^3] More info on https://www.smartvote.ch/en/wiki/methodology-smartspider/23_ch_nr?locale=en_CH. It contains annotations that allow for the visualization and deeper understanding of the positioning of parties according to predominant policy-domains in the political spectrum. We draw from the documentation provided by SmartVote where eight categories (considered stances on policy domains) are defined: open foreign policy, liberal economic policy, restrictive financial policy, law and order, restrictive migration policy, expanded environmental protection, expanded social welfare state, and liberal society. These categories are based on domains identified in Swiss political spectrum Hermann and Leuthold (2001, 2003), but that are generalized across European countries, as evidenced by the similarity with domains analyzed in cross-European studies such as the Chapel Hill Survey Jolly et al. (2022).
When answering ‘agree’ to a statement emphasizes any of the eight given policies, the statement is marked as a ‘agree’ with that policy domain, while disagreements with a policy are annotated as ‘disagree’. Three annotators with background in traditional or computational political science extended the annotations to the other countries. Table 1 shows that inter-annotator agreement – which is calculated with agreement between ‘agree’, ‘agree’ and no label per statement – is good. The final gold annotations are drawn from the majority votes. Note that some statements do not fall into any category. Therefore, the gold annotations contain 193 statements in total (tables 9 and 10, Appendix A provide examples and details).
4.3 Robustness to Statement Variations
We introduce three variants of each policy statement to test the models’ reliability (cf. statement variants, Figure 1 and robustness to statement variations in Figure 2).
@@ 4.3 Robustness to Statement Variations
The other type, which we call semantic opposite and denote $\text{Opp}(S)$, is a statement that takes the semantically opposite sense to the original one while not using an overt negation marker. A minimal number of words is modified to convert the semantic meaning of the sentence.
Each statement in the source language is annotated by a native speaker. Annotators are asked to create $\text{Neg}(S)$ by adding a marker corresponding to ‘not’ or ‘don’t’ in the source languages. As for $\text{Opp}(S)$, annotators are instructed to first try modifying the head verb in the statement or, if this is not possible, the focal adjective. If neither can be altered, they are asked to apply the minimal change necessary to invert the sentence’s meaning.
@@ 5.1 Models
Given that we formulate our prompts as zero-shot instruction, we opt for the instruction-following model families that vary in parameter size and contain different sets of pretraining data  [^4] We also evaluated the base models, but they do not produce any reliable answers, and are therefore not reported here.. Specifically, we focus on all size variants of LLama-2-Chat (7B, 13B, 70B) Touvron et al. (2023), the XXL variant of Flan-T5 (11B, Chung et al., 2022), Mistral Instruct (7B, Jiang et al., 2023) for open-sourced models and GPT-3.5 as a closed-source model. The models form three natural classes: small ($<$10B parameters, mistral-7b and llama2-7b), medium (between 10 and 19B (flanT5-xxl-11b and llama2-13b) and big ($>$20B, gpt3.5-20b and llama2-70b). All models utilize the top-p nucleus decoding approach, suggested by Holtzman et al. (2020), when generating responses, aiming to capture the model’s stance distribution through the sampling of the output (Cf. Appendix B for information on the implementation).
@@ 5.2 Prompt Design
When designing model inputs, we aim at creating templates that reliably elicit responses from models that can be clearly aligned with a defined stance, so either positive or negative. [^5] An example of an invalid response is I don’t know or I don’t have personal opinions. Considering recent research findings indicating that the meaning of prompts can be overshadowed by the choice of the target words as discussed in § 2, we aim at diversifying the choice for prompt templates from various angles. We distinguish between personal and impersonal templates (cf. prompt variants as shown in Figure 1). Personal templates formulate the instruction as a direct question (e.g. Do you agree or disagree?, Are you pro or con?), whereas impersonal are framed as objective tasks (Analyze the following statement into the labels “favorable” or “detrimental”, Classify the following statement as…). The context for evaluating the prompts is specified as Consider the long-term societal impact…. Additionally, we vary the wording of the stance (e.g. favorable, detrimental, advantageous, disadvantageous, support, oppose) to explore potential model biases in responding to specific wordings (cf. semantic label order, Figure 1). After a pilot experiment to test which prompts elicit most valid responses, we selected 3 personal and 3 impersonal prompt templates among 8 impersonal and 6 personal templates (Appendix B.1 details the selection process). Refer to the implemented prompt templates in Table 7, Appendix A.
@@
In order to test sensitivity of the models to subtle template changes each template is furthermore presented in two versions: the original one and the version where the order of the labels is swapped, e.g., if a template states, Analyze the following statement into the labels “favorable or detrimental”, the inverted-label version corresponds to “detrimental or favorable”. A reliable model is expected to yield the same response independent of label order (cf. robustness to label-order inversion, Figure 2).
@@
In addition to altering the statements, we modify the templates to investigate if the model maintains consistent stances with semantically equivalent templates. Previous research has demonstrated the impact of template variation on the results (Min et al., 2022; Khashabi et al., 2022). We hypothesize that variations in templates are likely to be an influential factor in shifts in the models’ generated stance.